OpenStudy (anonymous):

solve 3x^2-5x-11=0 find the x intersepts of f(x)=3x^2-5x-11.

hero (hero):

It's not complex

hero (hero):

\[D = b^2 - 4ac = (-5)^2 - 4(3)(-11) = 25 + 132 = 157\]

hero (hero):

\[x = \frac{\pm \sqrt{157} + 5}{6}\]

OpenStudy (anonymous):

any method ?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@Hero that's the answer i got,dat's why i called it a complex root

OpenStudy (anonymous):

@krypton : square root of a negative number gives you a complex number solution

myininaya (myininaya):

any number is complex

myininaya (myininaya):

4 is complex 4i is complex 4+4i is complex

OpenStudy (anonymous):

4+0i

OpenStudy (anonymous):

why do we have real and complex sets then ?

myininaya (myininaya):

real numbers is a subset of the complex imaginary numbers is a subset of the complex

myininaya (myininaya):

all numbers are complex

OpenStudy (anonymous):

If a complex number has b = 0, then it is a real number and it lies on the real axis in the complex plane

myininaya (myininaya):

Definition of Complex: \[Complex =\{a+bi | a \in \mathbb{R} , b \in \mathbb{R}\}\]

hero (hero):

Define complex in words only without using mathematical notation.

hero (hero):

What a meant to say is that the discriminant yields a real number, not a complex number. The definition of a complex number is a + bi

OpenStudy (anonymous):

The real line can be thought of as a part of the complex plane, and correspondingly, complex numbers include real numbers as a special case.

OpenStudy (anonymous):

... Special Case

hero (hero):

That's good to know. I'm not the one confused about it.

myininaya (myininaya):

Excuse me?

myininaya (myininaya):

I'm not confused either.

hero (hero):

I know you're not myininaya. I never said you were :P

OpenStudy (anonymous):

def : A complex number is neither purely real nor purely imaginary, but rather is a composite of the two ! Happy everyone ?

myininaya (myininaya):

Ok I guess I just understood you. sorry hero. :)

myininaya (myininaya):

misunderstood you* lol

myininaya (myininaya):

real numbers are complex imaginary numbers are complex real numbers+imaginary numbers are still complex

hero (hero):

she avoided using the word "pure" in her definitions.

myininaya (myininaya):

http://en.wikipedia.org/wiki/Complex_number

hero (hero):

So myininaya, which numbers are purely complex?

myininaya (myininaya):

I don't have a definition for "purely" complex numbers as I never heard of any set of numbers being defined at such sorry

OpenStudy (anonymous):

check out this book : An imaginary tale The Story of √-1 by Paul J. Nahim ( Professor Emeritus of Electrical Engineering at the University of New Hampshire )

hero (hero):

Guess there's no such thing as a pure complex number since it is a "mixture" of pure real and pure imaginary.

OpenStudy (anonymous):

4 is real 4+0i is complex

myininaya (myininaya):

4=4+0i

myininaya (myininaya):

4/2 is an integer just like 2 is

myininaya (myininaya):

because 4/2=2

myininaya (myininaya):

@jamesj @zarkon

hero (hero):

Why are you two reciting stuff you already know? Because I already know this stuff

OpenStudy (anonymous):